# Forming Auxiliary High-confident Instance-level Loss to Promote Learning from Label Proportions

### v1

@@ 3.2.2 Instance-level Adaptive Weight
At the instance level, the predicted distribution is derived from the prediction results of individual instances. Specifically, for each instance $\mathbf{x}_{i,j}$, we define $\mathbf{h}_{i,j}^{\mathrm{i}}=\left[\hat{y}^{\mathrm{w}}_{i,j,c}\right]_{c=1}^{C}$ to measure the confidence of the instance.
For the target distribution, we consider a label to have high confidence when its predicted probabilities are concentrated on a single class. Therefore, we define the target distribution as $\mathbf{\tilde{h}}_{i,j}^{\mathrm{i}}$, where only one class has a probability of 1, and the probabilities of all other classes are set to 0, i.e.,
$\mathbf{\tilde{h}}^{\mathrm{i}}_{i,j}$ $=\big[1,\underbrace{\vphantom{0}0,\ldots,0}_{C-1}\big].$ (9)
It is straightforward to calculate that the entropy of the target distribution is 0, i.e., $\mathcal{H}(\mathbf{\tilde{h}}^{\mathrm{i}}_{i,j})=0.$ With the same mapping function as the bag level, the instance-level adaptive weight is obtained with a hyper parameter $\beta_{\mathrm{i}}>0$:
$\omega^{\mathrm{i}}_{i,j}$ $=\sigma\left(\mathcal{H}(\mathbf{h}^{\mathrm{i}}_{i,j})-\mathcal{H}(\mathbf{\tilde{h}}^{\mathrm{i}}_{i,j});\beta_{\mathrm{i}}\right)$ (10)
$=\exp\left\{-\frac{\left[\mathcal{H}(\mathbf{h}^{\mathrm{i}}_{i,j})-\mathcal{H}(\mathbf{\tilde{h}}^{\mathrm{i}}_{i,j})\right]^{2}}{\beta_{\mathrm{i}}}\right\}$
$=\exp\left[-\frac{\mathcal{H}(\mathbf{h}_{i,j}^{\mathrm{i}})^{2}}{\beta_{\mathrm{i}}}\right].$
We provide a pseudo-code for one epoch of training for L2p-ahil, as shown in Algorithm 1.
Algorithm 1 Training process of L2p-ahil (one epoch)
Input: The LLP training dataset $\mathcal{D}$, weakly augmented data $\mathbf{x}^{\mathrm{w}}$, strongly augmented data $\mathbf{x}^{\mathrm{s}}$, model network $f_{\mathbf{W}}(\cdot)$
1:for $iter=1,2,\dots$ do
2:  Sample a mini-batch $\{(\mathbf{B}_{i},\mathbf{p}_{i})\}_{i=1}^{M}$ from $\mathcal{D}$;
3:  Model predictions $\mathbf{\hat{y}}^{\mathrm{w}},\mathbf{\hat{y}}^{\mathrm{s}}=f_{\mathbf{W}}(\mathbf{x}^{\mathrm{w}}),f_{\mathbf{W}}(\mathbf{x}^{\mathrm{s}})$;
4:  for $i=1$ to $M$ do
5:   Compute $\mathbf{\bar{p}}_{i}$ based on Eq. 3;
6:   for $j=1$ to NN do
7:    Generate $\mathbf{\tilde{y}}_{i,j}$ based on Eq. 4;
8:    Let $c=\arg\max(\mathbf{\hat{y}}^{\mathbf{w}}_{i,j})$;
9:    Compute $\omega^{\mathrm{b}}_{i,c}$ based on Eq. 8;
10:    Compute $\omega^{\mathrm{i}}_{i,j}$ based on Eq. 10;
11:    The adaptive weight $\omega_{i,j}=\omega^{\mathrm{b}}_{i,c}\cdot\omega^{\mathrm{i}}_{i,j}$;
12:   end for
13:  end for
14:  $\mathcal{L}_{\rm b}(\mathbf{p},\mathbf{\hat{y}}^{\mathrm{w}})=\frac{1}{M}\sum_{i=1}^{M}\mathcal{H}(\mathbf{p}_{i},\mathbf{\bar{p}}_{i})$;
15:  $\mathcal{L}_{\rm{i}}(\mathbf{\tilde{y}},\mathbf{\hat{y}}^{\mathrm{s}})=\frac{1}{M\cdot N}\sum_{i=1}^{M}\sum_{j=1}^{N}\omega_{i,j}\mathcal{H}(\mathbf{\tilde{y}}_{i,j},\mathbf{\hat{y}}_{i,j}^{\rm s})$;
16:  $\mathcal{L}(\mathbf{W})=\mathcal{L}_{\rm b}(\mathbf{p},\mathbf{\hat{y}}^{\mathrm{w}})+\lambda\mathcal{L}_{\rm{i}}(\mathbf{\tilde{y}},\mathbf{\hat{y}}^{\mathrm{s}})$;
17:  Update $\mathbf{W}$ with SGD to minimize $\mathcal{L}$;
18:end for
@@ 4.1 Experimental Setup
Datasets
We assessed L2p-ahil on four benchmark datasets for LLP: Fashion-MNIST [43], SVHN [27], CIFAR-10, CIFAR-100 [15] and Mini-ImagNet [42]. Fashion-MNIST is a dataset consisting of 60,000 training and 10,000 test examples. Each example is a 28x28 gray-scale image, representing 10 categories of fashion items. The SVHN dataset is composed of 32x32 RGB images of digits, featuring 73,257 training examples and 26,032 testing examples, along with an additional 531,131 training samples that we do not utilize in our experiments. CIFAR-10 and CIFAR-100 datasets each contain 50,000 training and 10,000 testing examples, with every example being a 32x32 color natural image. They are categorized into 10 and 100 classes, respectively. MiniImageNet is a subset of the ImageNet dataset, consisting of 100 selected classes. Each class contains 80 images for training and 20 images for testing, with all images resized to 64x64.
@@
For each dataset, with a specified bag size $M$, we randomly select $M$ samples from the training set $\mathcal{D}$ to form a bag $\mathbf{B}_{i}$. The samples in distinct bags do not overlap. We compute the class proportion information within each bag to guide the training, but real labels are not used during training. Following prior studies, we choose $M$ from the set $\{16,32,64,128,256\}$. As every dataset contains an equal number of samples across classes, the bags created through this method exhibit relatively balanced proportions, leading to more challenging LLP tasks, as evidenced by [46].
Table 1: Classification accuracy (mean±std) on CIFAR-10, CIFAR-100, SVHN, and Fashion-MNIST for different bag sizes. The highest accuracy is highlighted in bold. The notation ”*” indicates that the results of FLMm originate from its paper[45]. The notation ”-” indicates that the LLPFC is not suitable for CIFAR-100 or that FLMm did not report results for SVHN and Fashion-MNIST.
@@ 4.2 Implementation Details
We compare L2p-ahil against several state-of-the-art LLP methods. A) LLPFC [48]: This method offers a theoretically grounded approach to LLP by reducing it to learning with label noise, employing the forward correction (FC) loss. B) DLLP [1]: Traditional LLP methods primarily rely on the model’s output predictions to estimate the proportions of bags. C) LLP-VAT [40]: This method is based on consistency regularization, where the same sample under different data augmentations should yield similar prediction results. D) ROT [7] uses optimal transport to generate pseudo-labels for instance-level training combined with bag-level proportion loss. To ensure fairness, we employed the same data augmentation methods used in our approach for instance-level training. E) FLMm [45]: This method entails pre-training with contrastive learning using a larger encoder, ResNet-50, followed by fine-tuning utilizing the FLMm approach. $\ast\$) Fully supervised: Our fully-supervised approach uses instance-level ground-truth labels for training.
@@ 4.3 Results and Analysis
We present results on four benchmark datasets in Table 1. We observe that L2p-ahil significantly outperforms the majority of baseline methods across all benchmark datasets and for every bag size. Compared with FLMm, which utilizes a stronger pre-trained encoder, our method’s performance is only lower on CIFAR-100 with a bag size of 256. In other bag sizes, our method significantly surpasses FLMm, particularly on CIFAR-100 with bag sizes of 16 and 32, where we exceed their performance by 12.49% and 11.71%, respectively. For methods other than FLMm, our results exhibit a more gradual decline with increasing bag sizes. For instance, as the bag size increases from 16 to 256 on CIFAR-10, L2p-ahil’s performance drops by only 3.08%, whereas other methods show a decline of over 40%. Compared to ROT, which also uses pseudo-label self-training, we consistently achieve significantly better results in all cases. More importantly, L2p-ahil achieves results comparable to those of fully supervised learning models. For example, at a bag size of 16, L2p-ahil’s performance on simpler datasets like Fashion-MNIST and SVHN matches that of fully supervised learning. On more complex datasets such as CIFAR-10 and CIFAR-100, there’s only a slight performance degradation, lower than 2%. For the more challenging MiniImageNet dataset, L2p-ahil achieves the highest accuracy across all bag sizes.
@@
Figure 4: Plot of adaptive weight during training progresses on different dataset for 16, 64, 128 bag size. (a)(b) The value of the adaptive weight on CIFAR-10 and CIFAR-100.
@@
Figure 5: T-SNE visualization of the image representation on CIFAR-10 with bagsize = 256. Different colors represent the corresponding classes.
@@
We visualize the image representations generated by the feature encoder using t-SNE [41] in Fig.5. Different colors correspond to different ground-truth class labels. We use the CIFAR-10 dataset with a bag size of 256 and compare the t-SNE embeddings of three classic methods. For the traditional method DLLP, training the model solely with the bag-level proportion loss does not lead to effective representation learning. Although ROT introduces an instance-level loss through pseudo labels, it still falls short of achieving sufficient representation learning. In contrast, L2p-ahil generates well-separated clusters and more distinct representations, demonstrating its superior ability to learn high-quality representations.
@@ 4.4 Ablation Study
In this section, we perform an ablation study to assess the influence of two central elements of L2p-ahil: bag-level and instance-level adaptive weight. Experiments were conducted on the CIFAR-10 and CIFAR-100 datasets, with results presented in Table 2. The first row shows that removing both adaptive weight. With smaller bag sizes and fewer classes, the pseudo-labels generated by the model were relatively accurate, and the classification performance did not decline significantly, as seen with bag sizes from 16 to 64 on CIFAR-10. However, for larger bag sizes, such as 256, accuracy plummeted to merely 51.90%. On the more complex CIFAR-100, with a greater number of classes, the model’s performance deteriorated substantially. On CIFAR-100, we observe that the model’s performance drops sharply between adjacent bag sizes. The large number of classes and the reduction in supervisory information make the model more prone to generating inaccurate pseudo-labels, a phenomenon also observed in other weakly supervised domains. Without adaptive weight, the model is heavily impacted by incorrect pseudo-labels, leading to a significant performance drop. The second and third rows show that the model’s performance decreases when we remove the instance-level and bag-level adaptive weight, respectively. This decline is more pronounced with larger bag sizes, indicating that using only one type of weight is insufficient for accurately assessing the confidence of pseudo-labels.
@@ 4.5 Parameter Sensitivity
In this paragraph, we investigate how $\beta_{\mathrm{b}}$ and $\beta_{\mathrm{i}}$, key hyper parameters in L2p-ahil, influence its performance. We conduct experiments on CIFAR-10 and CIFAR-100 with a bag size of 128(Fig.6). For the 100-class task CIFAR-100, the best result is achieved with $\beta_{\mathrm{b}}=5$ and $\beta_{\mathrm{i}}=5$. For the 10-class task, represented by CIFAR-10, the optimal result is obtained with $\beta_{\mathrm{b}}=1$ and $\beta_{\mathrm{i}}=1$. We selected our hyper parameters $\beta_{\mathrm{b}}$ and $\beta_{\mathrm{i}}$ based on these tests. When $\beta_{\mathrm{b}}$ and $\beta_{\mathrm{i}}$ are small, the weights are highly sensitive to changes in entropy, and some accurate pseudo-labels may be assigned low weights, preventing the model from fully utilizing them. For example, when $\beta_{\mathrm{b}}$ = 2 and $\beta_{\mathrm{i}}$ = 2, the accuracy on CIFAR-100 is only 64.81%. Conversely, when $\beta_{\mathrm{b}}$ and $\beta_{\mathrm{i}}$ are too large, many smooth but inaccurate pseudo-labels are given high weights, which hampers the training of the classifier. For instance, with $\beta_{\mathrm{b}}$ = 40 and $\beta_{\mathrm{i}}$ = 40, the model fails to converge on CIFAR-100.
@@
Figure 6: The plots of joint sensitivity tests for hyper parameters $\beta_{\rm{b}}$ and $\beta_{\rm{i}}$ on CIFAR-10 and CIFAR-100 with a bag size of 128.
@@ 6 Limitations
Broader Impacts
The paper focuses solely on the technical aspects of LLP algorithms. Therefore, this work can benefit a wide range of machine learning researchers. Also, we do not expect our efforts to have any negative consequences.

### v2

@@ 3.2.2 Instance-level Adaptive Weight
At the instance level, the predicted distribution is derived from the prediction results of individual instances. Specifically, for each instance $\mathbf{x}_{i,j}$, we define instance-level predicted distribution $\mathbf{\pi}_{i,j}^{\mathrm{i}}=\left[\hat{y}^{\mathrm{w}}_{i,j,c}\right]_{c=1}^{C}$ to measure the confidence of the instance.
For the instance-level reference distribution, we consider a label to have high confidence when its predicted probabilities are concentrated on a single class. Therefore, we define the instance-level reference distribution as $\mathbf{\tilde{\pi}}_{i,j}^{\mathrm{i}}$, where only one class has a probability of 1, and the probabilities of all other classes are set to 0, i.e.,
$\mathbf{\tilde{\pi}}^{\mathrm{i}}_{i,j}$ $=\big{[}1,\underbrace{\vphantom{0}0,\ldots,0}_{C-1}\big{]}.$ (9)
It is straightforward to calculate that the entropy of instance-level reference distribution is 00, i.e., $\mathcal{H}(\mathbf{\tilde{\pi}}^{\mathrm{i}}_{i,j})=0$. With the same mapping function as the bag level, the instance-level adaptive weight is obtained with a hyper parameter $\beta_{\mathrm{i}}>0$:
$\omega^{\mathrm{i}}_{i,j}$ $=\sigma\left(\mathcal{H}(\mathbf{\pi}^{\mathrm{i}}_{i,j})-\mathcal{H}(\mathbf{\tilde{\pi}}^{\mathrm{i}}_{i,j});\beta_{\mathrm{i}}\right)$ (10)
$=\exp\left\{-\frac{\left[\mathcal{H}(\mathbf{\pi}^{\mathrm{i}}_{i,j})-\mathcal{H}(\mathbf{\tilde{\pi}}^{\mathrm{i}}_{i,j})\right]^{2}}{\beta_{\mathrm{i}}}\right\}$
$=\exp\left[-\frac{\mathcal{H}(\mathbf{\pi}_{i,j}^{\mathrm{i}})^{2}}{\beta_{\mathrm{i}}}\right].$
@@ 4.1 Experimental Setup
Datasets.
We assessed L2p-ahil on five benchmark datasets for LLP: Fashion-MNIST [44], SVHN [30], CIFAR-10, CIFAR-100 [16] and Mini-ImagNet [43]. Fashion-MNIST is a dataset consisting of 60,000 training and 10,000 test examples. Each example is a 28x28 gray-scale image, representing 10 categories of fashion items. The SVHN dataset is composed of 32x32 RGB images of digits, featuring 73,257 training examples and 26,032 testing examples, along with an additional 531,131 training samples that we do not utilize in our experiments. CIFAR-10 and CIFAR-100 datasets each contain 50,000 training and 10,000 testing examples, with every example being a 32x32 color natural image. They are categorized into 10 and 100 classes, respectively. MiniImageNet is a subset of the ImageNet dataset, consisting of 100 selected classes. Each class contains 80 images for training and 20 images for testing, with all images resized to 64x64.
@@
For each dataset, with a specified bag size $M$, we randomly select $M$ samples from the training set $\mathcal{D}$ to form a bag $\mathbf{B}_{i}$. The samples in distinct bags do not overlap. We compute the class proportion information within each bag to guide the training, but real labels are not used during training. Following prior studies, we choose $M$ from the set {16, 32, 64, 128, 256}. As every dataset contains an equal number of samples across classes, the bags created through this method exhibit relatively balanced proportions, leading to more challenging LLP tasks, as evidenced by [46].
Table 1: Classification accuracy (mean±std) on CIFAR-10, CIFAR-100, SVHN, Fashion-MNIST and MiniImageNet for different bag sizes. The highest accuracy is highlighted in bold. The symbol “*” indicates results for FLMm are reproduced directly from its original publication [45], while the dash symbol “-” signifies either the inapplicability of LLPFC to CIFAR-100 or the absence of reported results for SVHN, Fashion-MNIST, and MiniImageNet in FLMm’s experiments.
@@ 4.2 Implementation Details
A) LLPFC [48]: This method offers a theoretically grounded approach to LLP by reducing it to learning with label noise, employing the forward correction (FC) loss. B) DLLP [1]: Traditional LLP methods primarily rely on the model’s output predictions to estimate the proportions of bags. C) LLP-VAT [41]: This method is based on consistency regularization, where the same sample under different data augmentations should yield similar prediction results. D) ROT [9]: This method uses optimal transport to generate pseudo-labels for instance-level training combined with bag-level proportion loss. To ensure fairness, we employed the same data augmentation methods used in our approach for instance-level training. E) SoftMatch [6]: This method is one of the most important adaptive weighting techniques in weakly-supervised learning. We adopt the truncated Gaussian function, proposed in [6], as the instance-level adaptive weight. F) FLMm [45]: This method entails pre-training with contrastive learning using a larger encoder, ResNet-50, followed by fine-tuning utilizing the FLMm approach. $\ast$) Fully supervised: Our fully-supervised approach uses instance-level ground-truth labels for training.
@@ 4.3 Results and Analysis
We present results on five benchmark datasets in Table 1. We observe that L2p-ahil significantly outperforms the majority of baseline methods across all benchmark datasets and for every bag size. Compared with FLMm, which utilizes a stronger pre-trained encoder, our method’s performance is only lower on CIFAR-100 with a bag size of 256. In other bag sizes, our method significantly surpasses FLMm, particularly on CIFAR-100 with bag sizes of 16 and 32, where we exceed their performance by 12.49% and 11.71%, respectively. For methods other than FLMm, our results exhibit a more gradual decline with increasing bag sizes. For instance, as the bag size increases from 16 to 256 on CIFAR-10, L2p-ahil’s performance drops by only 2.08%, whereas other methods show a decline of over 40%. Compared to ROT, which also uses pseudo-label self-training, we consistently achieve significantly better results in all cases. SoftMatch performs well in certain cases, particularly when the bag size is small, such as 16 on CIFAR-10/100. However, as the bag size increases, its performance drops significantly. For example, when the bag size is 128 on CIFAR-100, the accuracy falls to only 2.12%. Although SoftMatch uses a similar technique to our method, the adaptive weighting it employs fails to address the issue of over-smoothing. More importantly, L2p-ahil achieves results comparable to those of fully supervised learning models. For example, at a bag size of 16, L2p-ahil’s performance on simpler datasets like Fashion-MNIST and SVHN matches that of fully supervised learning. On more complex datasets such as CIFAR-10 and CIFAR-100, there’s only a slight performance degradation, lower than 2%. For the more challenging MiniImageNet dataset, L2p-ahil achieves the highest accuracy across all bag sizes.
@@
Figure 4: Evolution of adaptive weights across training epochs for bag sizes 16, 64 and 256 on CIFAR-10 and CIFAR-100.
@@
Figure 5: t-SNE visualization of learned feature spaces on CIFAR-10 with bag size 256. Distinct colors correspond to different classes, where (a) DLLP and (b) ROT baselines exhibit overlapping clusters, while (c) our method L2p-ahil demonstrates clearer inter-class separation and tighter intra-class clustering, indicating more discriminative feature learning.
@@
We visualize the image representations generated by the feature encoder using t-SNE [42] in Fig. 5. Different colors correspond to different ground-truth class labels. We use the CIFAR-10 dataset with a bag size 256 and compare the t-SNE embeddings of three classic methods. For the traditional method DLLP, training the model solely with the bag-level proportion loss does not lead to effective representation learning. Although ROT introduces an instance-level loss through pseudo labels, it still falls short of achieving sufficient representation learning. In contrast, L2p-ahil generates well-separated clusters and more distinct representations, demonstrating its superior ability to learn high-quality representations.
@@ 4.4 Ablation Study
In this section, we perform an ablation study to assess the influence of two central elements of L2p-ahil: bag-level adaptive weight $\omega^{\rm b}$ and instance-level adaptive weight $\omega^{\rm i}$. Experiments were conducted on the CIFAR-10 and CIFAR-100 datasets, with results presented in Table 2. The first row shows that removing both adaptive weights. With smaller bag sizes and fewer classes, the pseudo-labels generated by the model were relatively accurate, and the classification performance did not decline significantly, as seen with bag sizes from 16 to 64 on CIFAR-10. However, for larger bag sizes, such as 256, accuracy plummeted to merely 51.90%. On the more complex CIFAR-100, with a greater number of classes, the model’s performance deteriorated substantially. On CIFAR-100, we observe that the model’s performance drops sharply between adjacent bag sizes. The large number of classes and the reduction in supervisory information make the model more prone to generating inaccurate pseudo-labels, a phenomenon also observed in other weakly supervised domains. Without adaptive weights, the model is heavily impacted by incorrect pseudo-labels, leading to a significant performance drop. The second and third rows indicate that the model’s accuracy decreases when the instance-level and bag-level adaptive weights are removed, respectively. This degradation intensifies with increasing bag sizes, highlighting that reliance on a single weight (bag-level or instance-level) inadequately captures pseudo-label confidence.
@@ 4.5 Parameter Sensitivity
In this section, we investigate how $\beta_{\mathrm{b}}$ and $\beta_{\mathrm{i}}$, two key hyper parameters in L2p-ahil that control the mapping functions at bag-level and instance-level respectively, influence model performance. We conduct experiments on CIFAR-10 and CIFAR-100 with a bag size 128 (Fig. 6). For the 100-class CIFAR-100 task, the best result is achieved with $\beta_{\mathrm{b}}=5$ and $\beta_{\mathrm{i}}=5$. For the 10-class task, represented by CIFAR-10, the optimal result is obtained with $\beta_{\mathrm{b}}=1$ and $\beta_{\mathrm{i}}=1$. We selected our hyper parameters $\beta_{\mathrm{b}}$ and $\beta_{\mathrm{i}}$ based on these tests. When $\beta_{\mathrm{b}}$ and $\beta_{\mathrm{i}}$ are small, the weights are highly sensitive to changes in entropy, and some accurate pseudo-labels may be assigned low weights, preventing the model from fully utilizing them. For example, when $\beta_{\mathrm{b}}$ = 2 and $\beta_{\mathrm{i}}$ = 2, the accuracy on CIFAR-100 is only 64.81%. Conversely, when $\beta_{\mathrm{b}}$ and $\beta_{\mathrm{i}}$ are too large, many smooth but inaccurate pseudo-labels receive high weights, which hampers classifier training. For instance, with $\beta_{\mathrm{b}}$ = 40 and $\beta_{\mathrm{i}}$ = 40, the model fails to converge on CIFAR-100.
@@
Figure 6: The plots of joint sensitivity tests for hyper parameters $\beta_{\rm{b}}$ and $\beta_{\rm{i}}$ on CIFAR-10 and CIFAR-100 with a bag size 128.